OpenStudy (anonymous):

Solve the systems by elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 Show all the steps

OpenStudy (anonymous):

I don't want someone to just give me the answer I need help figuring it out!

OpenStudy (wolfe8):

Would this help? http://www.themathpage.com/alg/simultaneous-equations-3.htm

OpenStudy (anonymous):

I tried to work the problem out using this method but I didn't get an answer that worked. :/

OpenStudy (wolfe8):

Can you show me everything you did step-by-step?

OpenStudy (anonymous):

i added _2x+2y+3z=0 to 2x+3y+3z=5 to get 5y+6z=5 then i added -2x-y+z=-3 to 2x+3y+3z=5 to get 2y+4z=2 and i am stuck from there

OpenStudy (wolfe8):

Try equation 3 minus equation 1 then add equation 2

OpenStudy (anonymous):

Add 2) and 3) \[-y + z + 3 y + 3 z = -3 + 5 \]\[2 y+4 z= 2 \]1) add 1) and 3)\[-2 x+2 y+3 z+2 x+3 y+3 z=0+5 \]\[5 y+6 z=5 \]You now have two equations in two unknowns:\[\{2 y+4 z=2,5 y+6 z=5\}\]@Danny_D Can you take it from here?

OpenStudy (anonymous):

that gives me x=5 right @wolfe8

OpenStudy (wolfe8):

Yeah it should give you x, then substitute x into another equation and do similar things.

OpenStudy (wolfe8):

My bad, I get x=-1

OpenStudy (anonymous):

i am so confused! how do we plug in x=-1 now?

OpenStudy (wolfe8):

Plug in x=-1 into all equations so now you only have simultaneous equation with 2 unknowns only.

OpenStudy (anonymous):

so -2(-1)+2y+3z=0

OpenStudy (wolfe8):

That's for the first equation. Do the same for the rest. Then you will have 2 unknowns left which should be simple.

OpenStudy (anonymous):

my brain is so dead right now, could you help me solve the simple equations too?

OpenStudy (wolfe8):

Well after that you will be left with equations 2y+3z=-2 -y+z=1 3y+3z+3 Add equation 1 to 2 times equation 2 should give you z.

OpenStudy (anonymous):

y+4z=3 times -y+z=1 right?

OpenStudy (wolfe8):

I meant 2 times equation 2, meaning equation 2 will become -2y+2z=2 Add that to equation 1 and get z.

OpenStudy (anonymous):

3z=0

OpenStudy (wolfe8):

If we did everything right, z is 0 then.

OpenStudy (anonymous):

ok then we can go through and put -2x+3=-2 -x+3=1 and 3x+3=3

OpenStudy (anonymous):

which makes x=0 as well

OpenStudy (wolfe8):

We already have x from the first step, remember? Now use z=0 in equation 3 that we get after putting in x.

OpenStudy (wolfe8):

That should give you y.

OpenStudy (anonymous):

thats what i ment x=0 but then i got three different answers for y when i plugged in these numbers!

OpenStudy (wolfe8):

Then you might want to check our workings. Sorry I can't myself. I'm studying for a test. But what we did is basically the procedure.

OpenStudy (anonymous):

thanks for your time

OpenStudy (wolfe8):

Or you could try robtobey's method up there.